OpenStudy (rose4825):

What is the approximate distance between the points (–2, 3, –5) and (2, –7, 4)? 11 14 17 23 My answer: 14

OpenStudy (rose4825):

@jim_thompson5910

OpenStudy (astrophysics):

Sounds good

jimthompson5910 (jim_thompson5910):

\[\Large d = \sqrt{\left(x_2-x_1\right)^2+\left(y_2-y_1\right)^2+\left(z_2-z_1\right)^2}\] \[\Large d = \sqrt{\left(2-(-2)\right)^2+\left(-7-3\right)^2+\left(4-(-5)\right)^2}\] \[\Large d = \sqrt{\left(2+2\right)^2+\left(-7-3\right)^2+\left(4+5\right)^2}\] \[\Large d = \sqrt{\left(4\right)^2+\left(-10\right)^2+\left(9\right)^2}\] \[\Large d = \sqrt{16+100+81}\] \[\Large d = \sqrt{197}\] \[\Large d \approx 14.0356688476181\] \[\Large d \approx 14\] So you have the correct answer. Nice job